40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1621     Entry time: Fri May 22 17:03:14 2009
Author: rob, steve 
Type: Update 
Category: PSL 
Subject: MOPA takes a holiday 

The MOPA is taking the long weekend off.

Steve went out to wipe off the condensation inside the MOPA and found beads of water inside the NPRO box, perilously close to the PCB board.  He then measured the water temperature at the chiller head, which is 6C.  We decided to "reboot" the MOPA/chiller combo, on the off chance that would get things synced up.  Upon turning off the MOPA, the neslab chiller display immediately started displaying the correct temperature--about 6C.  The 22C number must come from the MOPA controller.  We thus tentatively narrowed down the possible space of problems to: broken MOPA controller and/or clog in the cooling line going to the power amplifier.  We decided to leave the MOPA off for the weekend, and start plumbing on Tuesday.  It is of course possible that the controller is the problem, but we think leaving the laser off over the weekend is the best course of action.

 

 

ELOG V3.1.3-